I have to Evaluate this problem (g-6h)(g+6h)
I got =(g)^2-6h^2
= g^2-36h
is this correct?

Answer by venugopalramana(3286)

I have to Evaluate this problem (g-6h)(g+6h)
I got =(g)^2-6h^2
= g^2-36h
is this correct?
SEE BELOW MY COMMENTS
I have to Evaluate this problem (g-6h)(g+6h)
PUT BRACKETS TO AVOID MISTAKES I got =(g)^2-(6h)^2
SO YOU GET = g^2-36h^2
is this correct?ALMOST...YOU ARE GOOD...YOU CAN IMPROVE FURTHER..DEVELOP GOOD PRACTICES LIKE PUTTING APPROPRIATE BRACKETS...LITTLE MORE OF THEM IS BETTER THAN LESS AS YOUR PRESENT SMALL MISTAKE SHOWS...